OpenStudy (anonymous):

Help please! Use basic identities to simplify the expression sin^2 theta+ tan^2 theta + cos^2 theta

OpenStudy (saifoo.khan):

Go with cos.

OpenStudy (anonymous):

As the answer?

OpenStudy (saifoo.khan):

No. I mean simplify stuff in terms of cos

OpenStudy (anonymous):

@saifoo.khan

OpenStudy (saifoo.khan):

Yo?

OpenStudy (anonymous):

What did you get when you solved it? @saifoo.khan

OpenStudy (saifoo.khan):

What do you get?

zepdrix (zepdrix):

\[\large \sin^2\theta+\tan^2\theta+\cos^2\theta\] Hmm I guess you want to take advantage of your most basic `Square Identity` first. \[\large \sin^2\theta+\cos^2\theta=1\]Remember that one? :) Addition is commutative, so see how we can move the cosine around to apply this identity?

OpenStudy (anonymous):

Ok I'm following please go on from there @zepdrix

zepdrix (zepdrix):

So if we move the cosine over, since it's just addition, we get:\[\large \color{royalblue}{\sin^2\theta+\cos^2\theta}+\tan^2\theta\] So what does that simplify down to as our first step? :o

OpenStudy (anonymous):

@saifoo.khan I got sec^2?